OpenStudy (anonymous):

what is the slope of (3,3) and (-1,-1)

OpenStudy (radar):

\[m =\Delta y/\Delta x\]

OpenStudy (anonymous):

hey, let x1=3 and y1=3 x2= -1 and y2 = -1 then use the formula, slope(m) = y2-y1/x2-x1= 1

OpenStudy (anonymous):

use the formula y=mx + b after poulami's formula

OpenStudy (anonymous):

so what is the answer?

OpenStudy (radar):

m=(3-(-1))/3-(-1))=4/4=1

OpenStudy (radar):

Slope is 1

OpenStudy (anonymous):

what is the slope for (-3,1) and (1,5)

OpenStudy (radar):

Work it the same way. What are your two y values?

OpenStudy (anonymous):

1,5

OpenStudy (radar):

right, what is the difference between those points, 5-1=???

OpenStudy (anonymous):

5-1=4

OpenStudy (radar):

Yes, 4 is the numerator. Now work out the denominator. What are the x values?

OpenStudy (anonymous):

-3-1=-2

OpenStudy (radar):

Almost. The difference between -3 and +1 if you were at -3 how many units would you have to move to get to +1. We are trying to get the distance from -3 to +1

OpenStudy (radar):

-3 - 1=-4

OpenStudy (anonymous):

so the answer is 4/-4

OpenStudy (radar):

1-(-3)=1+3=4 We have to go the same direction we went with the y value

OpenStudy (radar):

so the answer is 4/4 or 1 again.

OpenStudy (anonymous):

the slope for (-1,-1) and(3,3) is the slope 2 and 4

OpenStudy (radar):

What you must do is to label your two points (in your mind) as P1 and P2\[m=(y _{2}-y _{1})/(x _{2}-x _{1})\]

OpenStudy (radar):

Which point do you want to call P1?

OpenStudy (radar):

O.K. I will make the assignment (-1,-1) is P1, then (3, 3) is P2 then slope m=(3-(-1))/(3-(-1))=4/4=1 slope is 1

OpenStudy (anonymous):

(-3,1) (1,5) (3,3) (-1,-1) all these sides are each what